OpenStudy (anonymous):

Solve the exponential equation. Use a calculator to obtain a decimal approximation, correct to two decimal places, for the solution. e x + 5 = 2

OpenStudy (anonymous):

\[e ^{x+5} = 2\]

OpenStudy (anonymous):

Take natural log on both sides \[x + 5 = \ln 2\]\[x = \ln 2 - 5\]

OpenStudy (anonymous):

-4.31

OpenStudy (anonymous):

lol

OpenStudy (turingtest):

oh two DECIMAL places.. -4.31

OpenStudy (anonymous):

Wooo 3rd time the charm xD

OpenStudy (turingtest):

this server sometimes... :/

OpenStudy (turingtest):

hooray...